2) You are asked the conduct the following hypothesis test concerning a population proportion: H0: prop...

2)

You are asked the conduct the following hypothesis test concerning a population proportion:

H0: prop = 0.43

HA: prop > 0.43

You draw a sample of size 437 from the population. There are 198 individuals in the sample who satisfy the criteria in which we are interested. Calculate the p-value for this test.

Round your answer to 4 decimal places.
Answer #1

2)

Solution :

This is the right tailed test .

= x / n = 198 / 437 = 0.4531

Test statistic = z

= - P0 / [P0 * (1 - P0 ) / n]

= 0.4531 - 0.43 / [(0.43 * 0.57) / 437]

= 0.975

P-value = 0.1648